OpenStudy (anonymous):

write an equation in point slope form of the line that passes through the points (-6,-3) and (6,-4)

OpenStudy (dls):

Find the slope, you know how to?

OpenStudy (dumbcow):

here is point slope form: \[y-y_1 = m(x-x_1)\] where (x1,y1) is a point and m is the slope \[m = \frac{y_2 -y_1}{x_2 -x_1}\]